# PyLO: Towards Accessible Learned Optimizers in PyTorch

### v1

@@ 5 Demonstrating the real world use of PyLO
Figure 4: Weight decay and learning rate schedule (LRS) ablation. In PyLO, it is easy to equip VeLO and $\mu$LOM with weight decay and learning rate schedules. We sweep different values and find that in some cases learned optimizers improve.
@@ 5.1 Training Vision Transformer on ImageNet-1K
Experimental Configuration: We train ViT-Base/16 models (86M parameters) for 480 epochs (150k steps) on ImageNet-1K following established protocols (Wightman, 2019). We apply standard augmentation techniques (RandAugment, CutMix, Mixup) to ensure realistic training conditions that reflect practical deployment scenarios. We employ a batch size of 4096 and conduct these experiments on Nvidia H100 GPUs.
Results As shown in Table 2 VeLO demonstrates competitive performance in this context, achieving 78.45% top-1 accuracy compared to AdamW’s 77.22%. This scenario demonstrates a case where the learned optimizer outperforms Adam. $\mu$LO exhibits training divergence after 65,000 steps due to its limited 5,000-step meta-training horizon, achieving only 62.14% peak validation accuracy.
@@ 5.2 Language model pre-training on FineWeb
Experimental Configuration We pre-train a $355$M parameter GPT-2 style transformers on FineWeb-EDU (Lozhkov et al., 2024) for $10$B tokens, employing a batch size of $512$ and sequences of length $1024$ (see section G for additional hyperparameter details).
Results Table 2 reports the final loss achieved for all optimizers in our study, while figure 7 of the appendix plots the training curves. We observe that VeLO outperforms AdamW and $\mu$LOM, reaching the lowest final loss. This demonstrates that PyLO has the potential to deliver optimization performance improvements to PyTorch users.
@@ 5.3 Combining scheduler and decoupled weight decay
Effect of Learning Rate Scheduler We equip $\mu$LOM and VeLO with a cosine annealing LRS and tune the maximum learning rate (MaxLR). Figure 4 reports ViT and GPT training and accuracy curves for the best-performing MaxLR while training curves for more values are reported in the appendix (F). We observe that $\mu$LOM benefits substantially from explicit scheduling, extending stable training from 65k to 150k steps and reaching 71% Top-1 accuracy on ImageNet and improving loss in language model pre-training. VeLO shows minimal scheduling improvements, suggesting different internal adaptation mechanisms.
Effect of Weight decay We equip $\mu$LOM and VeLO with decoupled weight decay and tune the decay coefficient, $\lambda$. Figure 4 reports ViT and GPT training and accuracy curves for the best-performing $\lambda$ while training curves for more values are reported in the appendix (F). We observe that $\mu$LOM benefits from weight decay for training GPT but not VeLO. Neither optimizer benefits from weight decay for ViT tasks.
@@ 6 Conclusion
In conclusion, we have introduced PyLO, an open-source repository for state-of-the-art learned optimizers in PyTorch. PyLO breaks down the barriers to learned optimizers adoption by integrating directly with torch.optim.Optimizer and the HuggingFace ecosystem, allowing users to directly leverage learned optimizers in their existing code and easily download and share learned optimizer weights. For the best performance, PyLO implements a CUDA accelerated forward pass for the small_fc_lopt architecture, which we demonstrate has superior scaling properties and lower memory overhead than the original JAX implementation. When benchmarking the PyLO implementation of $\mu$LOM and VeLO against AdamW, we observed that VeLO can outperform AdamW on transformer language modeling and ViT tasks. We also ran ablations facilitated by PyLO’s PyTorch integration and found that learning rate schedules and weight decay can substantially improve the performance of certain learned optimizers.
Acknowledgments
We acknowledge support from FRQNT New Scholar [E.B.] and the FRQNT Doctoral (B2X) scholarship [B.T., A.M.]. We also acknowledge resources provided by Compute Canada, Calcul Québec, and Mila.
Appendix A Comparison to Existing Learned Optimization Frameworks
Repository Overviews
Open-L2O (Chen et al., 2022) offers a comprehensive benchmarking suite designed primarily for evaluating learned optimizers (L2O) across a range of optimization problems. It includes both model-based and model-free L2O methods, facilitating reproducibility and fair comparisons but does not focus on practical deployment for real-world tasks.
Google’s learned_optimization (Metz et al., 2022a) is a research-oriented repository that provides a robust JAX-based framework for training, designing, evaluating, and applying learned optimizers. It extensively supports meta-training mechanisms, including Evolution Strategies (ES) and Predictive ES (PES), but lacks user-friendly interfaces for applying these optimizers in practical, production-oriented settings.
PyLO (ours) addresses critical barriers that have limited the adoption of learned optimizers in the broader machine learning community by providing an accessible, performance-focused PyTorch implementation. PyLO features seamless integration with PyTorch workflows, CUDA acceleration for computational efficiency, and HuggingFace ecosystem integration for easy sharing of optimizer weights and community-driven development.
Comparative Analysis
PyLO distinguishes itself from Open-L2O and Google’s learned_optimization in several key aspects:
  1. 1.
    Framework Compatibility: PyLO is built entirely in PyTorch, ensuring compatibility with the predominant framework used by practitioners, unlike Google’s learned_optimization, which is JAX-based.
  2. 2.
    Decoupled Meta-testing: PyLO exclusively provides optimized implementations for meta-testing without the complexity and overhead associated with meta-training code, unlike the other repositories.
  3. 3.
    HuggingFace Integration: PyLO integrates deeply with the HuggingFace Hub, facilitating standardized sharing, versioning, and benchmarking of task-specific optimizers, a capability not provided by the other repositories.
  4. 4.
    Computational Efficiency: By leveraging specialized CUDA kernels, PyLO significantly reduces optimizer step overhead, enabling practical deployment for large-scale models, a substantial improvement over naive implementations.
The strengths of PyLO position it uniquely as both researcher- and practitioner-friendly, enabling efficient exploration and practical use of state-of-the-art learned optimization techniques.
Table 3: Comparison of Learned Optimization Repositories
Repository HF PyTorch Decoupled Paper Github
Open-L2O (Chen et al., 2022) Open-L2O
Learned Optimization (Metz et al., 2022a) learned_optimization
PyLO (ours) In submission PyLO
Appendix B Extended challenges of L2O adoption for the Wider ML community
This section examines the critical barriers hindering the widespread adoption of learned optimizers within the broader machine-learning community. Despite their demonstrated potential for reduced training wall-clock time across various domains(Metz et al., 2022a; Chen et al., 2022), several structural and technical challenges have limited the integration of learned optimizers into standard machine learning workflows. The main barriers can be summarized as follows:
Several critical barriers have hindered the more widespread adoption of learned optimizers for deep learning tasks in our community: 1) overfocus on meta-learning in current frameworks, 2) absence of a PyTorch implementation for the most performant learned optimizers available, 3) the inability to effectively share optimizer weights, 4) the absence of learned optimizer benchmarking for popular machine learning tasks, and 5) the optimizer step overhead of learned optimizers.
  1. 1.
    The overfocus on meta-learning in current frameworks.
  2. 2.
    The absence of a PyTorch implementation for the most performant learned optimizers available.
  3. 3.
    The lack of a standardized way to effectively share learned optimizer weights.
  4. 4.
    The absence of learned optimizer benchmarking for popular machine learning tasks.
  5. 5.
    The optimizer step overhead of learned optimizers.
Overfocus on meta-learning
A primary obstacle to L2O adoption stems from the disparity between current learned optimization frameworks and established training paradigms. Existing libraries, such as the JAX (Bradbury et al., 2018)-based learned optimization framework (Metz et al., 2022a, b), prioritize meta-training capabilities over practical deployment. While these frameworks do provide mechanisms for utilizing pre-trained learned optimizers, the interfaces diverge significantly from those familiar to practitioners (Wightman, 2019; Wolf et al., 2020; von Platen et al., 2022). The evaluation and application of learned optimizers appear as secondary objectives, with insufficient emphasis on seamless integration with conventional training pipelines. This architectural choice highlights the necessity of decoupling meta-learning and meta-testing phases, particularly for researchers and practitioners who seek to leverage pre-trained optimization algorithms without engaging in the complexities of meta-training.
Pytorch v.s. JAX
A second significant barrier relates to the underlying deep learning frameworks. While JAX (Bradbury et al., 2018) offers powerful functional programming capabilities and high flexibility for research workflows, it lacks the mature ecosystem and widespread adoption of more established frameworks. In contrast, PyTorch (Paszke et al., 2019) has developed into the dominant deep learning framework with extensive community support (He, 2019) and a rich ecosystem of pre-trained models and resources (Wightman, 2019; von Platen et al., 2022; Wolf et al., 2020). This prevalence stems partly from PyTorch’s object-oriented architecture and modular design principles, which facilitate extension, ease-of-use, and customization. The absence of a comprehensive learned optimizer library that conforms to PyTorch’s optimizer interface presents a substantial obstacle, as it necessitates significant codebase modifications for implementation and experimentation. Consequently, there exists a pressing need for a modular L2O framework that seamlessly integrates with PyTorch’s optimizer class hierarchy. Such integration, particularly when combined with standardized weight serialization and distribution through established repositories, would substantially accelerate both the circulation and scaling of learned optimization techniques.
Absence of standardized benchmarking
A third critical limitation concerns evaluation methodologies and benchmarking procedures. Current L2O framework Metz et al. (2022a) , typically evaluate optimization algorithms across an extensive but potentially misaligned task distribution. These evaluations predominantly focus on training loss trajectories rather than generalization performance metrics that are of primary interest to practitioners. Moreover, recent works like VeLO(Metz et al., 2022b) omit crucial experimental details, including baseline hyperparameter configurations , impeding reproducibility and comparative analysis(Rezk et al., 2023). This deficiency underscores the necessity for standardized evaluation frameworks that enable systematic comparison of conventional and learned optimizers on practically relevant tasks. Such frameworks should prioritize contemporary challenges in vision and language model pre-training rather than theoretical convex optimization problems that, while analytically tractable, often fail to capture the complexities of modern neural network training dynamics.
Sharing learned optimizer weights
Pre-trained models for language understanding (DeepSeek-AI et al., 2025; Touvron et al., 2023), image classification (Dosovitskiy et al., 2020; Radford et al., 2021), semantic segmentation (Kirillov et al., 2023), and a number of other tasks are routinely shared via platforms such as HuggingFace Hub (HuggingFace, ), enabling practitioners to leverage computational investments made by well-resourced institutions. This collaborative open-model ecosystem has accelerated innovation and reduced redundancy in these domains through its expansion of access to state-of-the-art models. Despite this revolution in model sharing, we lack seamless sharing in learned optimizer weights. We suggest that foundational Learned Optimizer (LO) models can and should be integrated into the community following similar distribution mechanisms (e.g., HuggingFace) to those established for neural network weights. Our objective is to facilitate seamless integration of learned optimizers into existing PyTorch (Paszke et al., 2019) workflows with minimal friction, addressing the significant gap in accessibility for the predominant deep learning framework used by practitioners. Furthermore, by establishing infrastructure for distributing optimizer weights, we aim to catalyze a parallel ecosystem where learned optimizer rules can be shared and deployed across diverse problem domains.
Per-step overhead of learned optimizers
A significant obstacle to the widespread adoption of learned optimizers is the computational overhead they potentially introduce to already resource-intensive training setups. Deep learning workflows operate under strict time and computational constraints, with practitioners demonstrating marked reluctance to try new optimization strategies that might increase training durations, regardless of potential convergence benefits. This challenge necessitates meticulous implementation of learned optimization algorithms to ensure competitive computational efficiency. Our approach addresses this critical constraint through custom CUDA implementations that substantially reduce optimizer step time compared to native PyTorch implementations. By implementing the learned optimizer update as a neural network pass directly in CUDA and strategically storing intermediate activations in registers, we minimize the need to write activations to global GPU memory. This optimization achieves substantial reductions in wall-clock time for optimization steps and render learned optimizers practically viable within existing computational budgets, removing a significant barrier to their adoption by the broader machine-learning community.
Appendix C Extended Code design
We implement a modular architecture partitioned into three main components to realize our project vision. This design philosophy prioritizes both usability for practitioners and extensibility for researchers developing novel learned optimizers.
CUDA Acceleration (pylo.csrc) The computational demands of learned optimizers present significant implementation challenges. Unlike traditional optimization algorithms that apply simple, closed-form update rules, learned optimizers require evaluating a small multilayer perceptron (MLP) for each parameter in the optimizee. This can introduce substantial computational overhead for large models, creating a critical bottleneck that limits practical deployment. Our analysis revealed that standard PyTorch neural network modules are inadequately optimized for this specific use case. These modules are primarily designed for batched operations on image or language data rather than the unique computational pattern of learned optimizers, where many small MLPs must be evaluated in parallel across millions of parameters.
Drawing inspiration from the architecture-aware optimizations (Müller et al., 2021; Dao et al., 2022) we implemented a specialized CUDA kernel for the small_fc_lopt learned optimizer. This implementation strategically leverages the GPU memory hierarchy to address the primary bottleneck: memory bandwidth rather than computational throughput. Our CUDA implementation employs two key optimization strategies: (1) Register-Based Computation and (2) On-the-Fly Feature Computation. The first utilizes GPU registers to store intermediate activations during forward propagation through the optimizer’s MLP. This approach minimizes high-latency global memory accesses by keeping frequently accessed data in the fastest available memory tier. The second avoids storing normalized features in global memory, recalculating them when needed, and trading redundant computation for reduced memory bandwidth.
Appendix D Extended Experimental Results
D.1 Benchmarking LO step times in PyLO
Model Optimizer Batch Size Samples per sec Step time (ms) Forward time (ms) Backward time (ms) Optimizer step time(ms) Num parameters(M) Vit-B-16 AdamW 32.00 524.09 60.55 17.52 38.13 4.90 86.57 Adafactor 425.51 74.69 18.99 small_fc_lopt (naive) 39.36 812.51 756.80 small_fc_lopt (CUDA) 205.59 155.16 99.59 Vit-S-16 AdamW 32.00 1,116.12 28.17 7.65 18.27 2.25 22.05 Adafactor 711.97 44.47 18.53 small_fc_lopt (naive) 109.03 293.01 266.90 small_fc_lopt (CUDA) 382.73 83.28 57.45 Vit-L-16 AdamW 32.00 175.95 181.36 52.18 113.89 15.28 304.33 Adafactor 156.64 203.34 37.23 small_fc_lopt (naive) 11.64 2,748.52 2,582.30 small_fc_lopt (CUDA) 70.69 451.69 285.80 GPT2-125 m AdamW 4.00 17.80 224.72 73.20 144.24 7.29 125.26 Adafactor 17.02 235.07 17.76 small_fc_lopt (naive) 3.38 1,182.33 964.89 small_fc_lopt (CUDA) 11.71 341.46 124.09 GPT2-355 m AdamW 4.00 6.56 609.83 197.41 392.29 20.12 355.92 Adafactor 6.40 624.64 35.11 small_fc_lopt (naive) 1.16 3,461.96 2,872.17 small_fc_lopt (CUDA) 4.40 908.69 319.14 GPT2-1B AdamW 4.00 2.88 1,387.51 442.49 896.81 48.21 912.95 Adafactor 2.87 1,395.03 54.08 small_fc_lopt (naive) OOM OOM OOM small_fc_lopt (CUDA) 1.98 2,025.26 681.78
Table 4: Breakdown of time spent per training step, averaged over 40 steps. The CUDA-based implementation reduces overhead from the learned optimizer by minimizing memory usage and optimizer time, ultimately achieving competitive throughput compared to traditional optimizers.
We extend our analysis beyond the baseline results to examine per-step computational overhead across varying model scales representative of contemporary pretraining workloads. Table 4 presents a comprehensive timing breakdown across forward pass, backward pass, and optimizer step durations for both vision and language model architectures. Traditional optimizers including AdamW (Loshchilov & Hutter, 2019) and Adafactor (Shazeer & Stern, 2018) exhibit minimal computational overhead relative to the forward and backward passes. Our CUDA kernel implementations demonstrate substantial reductions in learned optimizer overhead, with improvements that scale favorably with model size. Notably, for ViT-L-16, our small_fc_lopt(CUDA) implementation achieves a 9× reduction in optimizer step time compared to the naive implementation. This performance advantage becomes increasingly pronounced at scale. Consistent with our earlier findings regarding memory limitations—where JAX implementations encounter out-of-memory errors for 1B parameter language models—the naive learned optimizer implementation similarly fails at this scale. Our CUDA implementation addresses these scalability constraints while maintaining the computational efficiency gains observed with increasing batch sizes, making learned optimization viable for large-scale pretraining scenarios.
(a) Width
(b) Depth
Figure 5: Step time for deeper and wider MLPs. Increasing the width and depth of MLPs provides a systematic way to measure the impact of larger and more tensors on optimizer step time. Subfigure (a) reports optimizer step time as the hidden size of the 5-layer optimizee MLP is increased, while Subfigure (b) reports step time as the depth of a 256 hidden-layer MLP is increased. We observe that in each case, the CUDA implementation reduces the optimizer step time by more than an order of magnitude.
Scaling analysis with MLP: To systematically characterize the scaling behavior of our implementation, we conduct controlled experiments using synthetic MLP architectures with varying architectural parameters. Figure 5 presents two complementary scaling analyses that isolate the effects of network width and depth on optimizer computational overhead. Figure 5(a) examines the relationship between hidden layer dimensionality and optimization time for a fixed 5-layer MLP architecture. The results reveal a stark performance disparity: the naive implementation of small_fc_lopt exhibits optimization times an order of magnitude greater than our CUDA implementation and encounters memory limitations beyond hidden dimensions of 256. While our CUDA implementation remains slower than traditional optimizers (AdamW and Adafactor), it demonstrates favorable scaling characteristics that maintain computational tractability across the tested range.
Complementary analysis in Figure 5(b) evaluates depth scaling behavior using MLPs with fixed width of 256 units but varying layer counts (depth). The naive implementation consistently operates in a computationally prohibitive regime, requiring over 400ms for optimization steps, while our CUDA implementation maintains optimization times below 50ms—approaching the performance envelope of traditional optimizers despite the inherent complexity of learned optimization. These scaling experiments demonstrate that our CUDA implementation successfully addresses the computational bottlenecks that render naive learned optimizer implementation impractical for realistic model architectures.
@@ D.1 Benchmarking LO step times in PyLO
Figure 6: Vision Transformer training dynamics. Training loss curves for Vision Transformers trained on ImageNet-1k classification across three model configurations: (a) ViT-small with patch size 16, (b) ViT-base with patch size 16, and (c) ViT-large with patch size 16. We compare AdamW with cosine learning rate scheduling (blue), VeLO learned optimizer (orange), and $\mu$LO learned optimizer (black) over 150,000 training steps with batch size 4,096. $\mu$LO exhibits early training instability across all scales, while VeLO demonstrates competitive or superior convergence properties, particularly evident in the base model configuration where it achieves lower final loss than the AdamW baseline.
Appendix E Training Curves for the Experiments
This section presents a comprehensive empirical evaluation comparing learned optimizers VeLO (Metz et al., 2022b) and $\mu$LO (Thérien et al., 2024) against the standard AdamW optimizer (Loshchilov & Hutter, 2019; Kingma & Ba, 2017) with cosine learning rate scheduling (Loshchilov & Hutter, 2017). Our evaluation spans Vision Transformer architectures of varying scales and GPT-2 models, extending beyond the main paper’s validation accuracy results for ViT-Base/16 and final loss metrics for GPT-2 to provide detailed training dynamics across multiple configurations.
Vision Transformer Experiments: We evaluate three ViT configurations—Small, Base, and Large with patch size 16—on ImageNet-1k classification tasks. All configurations maintain consistent training protocols with 150,000 optimization steps and batch sizes of 4,096 across model variants, using hyperparameters detailed in Table 6. The extended evaluation presented in Figure 6 reveals critical scale-dependent optimization characteristics. MuLO demonstrates pronounced early divergence patterns across both small and large model configurations, reinforcing our hypothesis that its constrained meta-training horizon limits generalization to extended optimization trajectories. In contrast, VeLO exhibits remarkable consistency in convergence properties across the full spectrum of model scales examined, maintaining stable training dynamics while achieving competitive performance metrics relative to the AdamW baseline.
Language Modeling Experiments For GPT-2 pretraining on FineWeb-EDU data, we train models for 19,073 steps with batch size 512 and sequence length 1024. Complete hyperparameters are presented in Table 8, with full training curves shown in Figure 7. VeLO demonstrates competitive performance against AdamW with cosine scheduling, while MuLO exhibits significantly degraded performance across the extended training horizon, consistent with the stability patterns observed in vision tasks.
@@ Appendix E Training Curves for the Experiments
Appendix F Extended Ablation Studies
@@
Vision Transformer Ablations Figures 8 and 9 demonstrate the variation of training dynamics with varying maximum learning rates and weight decay values for the ViT-B/16 model. The results reveal that VeLO exhibits minimal sensitivity to weight decay and scheduler modifications in this experimental setup, suggesting robust inherent regularization properties. Conversely, MuLO shows improved training horizon stability when augmented with cosine scheduling, enabling successful completion of the full 150,000 training steps. Weight decay modifications yield marginal improvements for MuLO across the evaluated parameter ranges.
@@
The language modeling experiments, presented in Figures 10 and 11, demonstrate more pronounced sensitivity to weight decay and scheduling compared to vision tasks. VeLO shows measurable performance improvements when augmented with weight decay and scheduling for specific parameter configurations. Similarly, MuLO benefits substantially from both weight decay regularization and cosine scheduling in the language modeling domain, suggesting that learned optimizers may benefit from scheduling and weight decay.
@@
Figure 12 presents an additional ablation examining the impact of separate versus concatenated query-key-value (QKV) matrices on the optimizee’s training loss. We observe that separating the QKV matrices in attention blocks improves performance across VeLO and $\mu$LOM.
@@
(a) $\mu$LO Decoupled Weight decay ablation
(b) $\mu$LO Cosine annealing Ablation
Figure 9: Training ViT-B/16 with $\mu$LO on Imagenet1k with decoupled weight decay and cosine annealing. We see that the training horizon of $\mu$LO is improved with using a scheduler that helps it to run up to 150000 training steps. We also see that weight decay did not have substantial improvement
@@
(a) $\mu$LO Decoupled Weight decay ablation
(b) $\mu$LO Cosine annealing Ablation
(c) $\mu$LO LR Tuning Ablation
@@
(b) $\mu$LO
Figure 12: Separate QKV weight ablation. We ablate using concatenated or separate QKV matrices in attention layers. From the perspective of the learned optimizer, these two settings will be treated differently. We observe that performance improves for both $\mu$LOM and VeLO when QKV matrices are separated.
Appendix G Hyperparameters for Experiments

### v2

@@ 8 Demonstrating the real-world use of PyLO
[…]
@@ 8.1 Training Vision Transformer on ImageNet-1K
Experimental Configuration: We train ViT-Base/16 models (86M parameters) for 480 epochs (150k steps) on ImageNet-1K following established protocols Wightman (2019). We apply standard augmentation techniques (RandAugment, CutMix, Mixup) to ensure realistic training conditions that reflect practical deployment scenarios. We employ a batch size of 4096 and conduct these experiments on Nvidia H100 GPUs.
[…]
@@ 8.2 Language model pre-training on FineWeb
Experimental Configuration
We pre-train a 355355M parameter GPT-2 style transformers on FineWeb-EDU Lozhkov et al. (2024) for 1010B tokens, employing a batch size of $512$ and sequences of length $1024$ (see section D for additional hyperparameter details).r more details.
Results
[…]
@@ 8.3 Combining scheduler and decoupled weight decay
Effect of Learning Rate Scheduler
[…]
Effect of Weight decay
[…]
@@ 9 Conclusion
We have presented PyLO, an open-source PyTorch library that removes key barriers to the practical use of learned optimizers in large-scale deep learning. PyLO offers drop-in implementations of state-of-the-art methods such as VeLO and small_fc_lopt through the standard torch.optim.Optimizer interface, with seamless Hugging Face Hub integration for model sharing. Central to our systems contributions are custom CUDA kernels that fuse feature construction, normalization, and MLP inference, and a distributed implementation that distributes tensors across multiple devices. This achieves 86–88% lower overhead than naive PyTorch, and can efficiently scale to training a 1B-parameter transformer on a single A100 GPU. Our evaluation shows that PyLO reduces learned optimizer overhead to acceptable levels at scale and they can deliver competitive accuracy on ImageNet and large-scale language pretraining. As future work, we envision PyLO enabling (1) hardware-optimizer co-design (2) behavior-cloned distillation of expensive second-order optimizers such as SOAPVyas et al. (2025). PyLO thus establishes a robust foundation for community-driven research and practical deployment of learned optimizers in modern ML systems.
Appendix A Enhanced Performance on NVIDIA H100 GPUs
To further demonstrate the scalability of our approach on state-of-the-art hardware, we evaluate its performance on NVIDIA H100 GPUs, which offer advanced features. These accelerators are particularly beneficial for large matrix operations, such as those with dimensions $8192\times 8192$—corresponding to the hidden size of a 70B-parameter transformer model.
In our enhanced implementation (denoted small_fc_lopt-PyLO-CUDA++), we introduce the following optimizations:
  • Reduction in the number of global MLP reads to minimize HBM accesses;
  • Asynchronous memory copies to overlap data movement with computation; and
  • Full utilization of tensor cores and WGMMA instructions for high-throughput matrix multiplications on large tensors.
These modifications substantially reduce the optimizer step time, effectively closing the performance gap relative to state-of-the-art baselines for large-scale workloads. The results are summarized in Figure 9.
Figure 9: Optimizer step time comparison on NVIDIA H100 GPUs for $8192\times 8192$ matrices. Our enhanced PyLO-CUDA++ implementation (dark green) achieves a $2.1\times$ speedup over the baseline PyLO-CUDA (light green). Muon (pink) and Adam (orange) given for reference
Appendix B Training Curves for the Experiments
@@
[…]
@@
[…]
Vision Transformer Experiments: We evaluate three ViT configurations—Small, Base, and Large with patch size 16—on ImageNet-1k classification tasks. All configurations maintain consistent training protocols with 150,000 optimization steps and batch sizes of 4,096 across model variants, using hyperparameters detailed in Table 6. The extended evaluation presented in Figure 10 reveals critical scale-dependent optimization characteristics. MuLO demonstrates pronounced early divergence patterns across both small and large model configurations, reinforcing our hypothesis that its constrained meta-training horizon limits generalization to extended optimization trajectories. In contrast, VeLO exhibits remarkable consistency in convergence properties across the full spectrum of model scales examined, maintaining stable training dynamics while achieving competitive performance metrics relative to the Adam baseline.
Language Modeling Experiments For GPT-2 pretraining on FineWeb-EDU data, we train models for 19,073 steps with batch size 512 and sequence length 1024. Complete hyperparameters are presented in Table 8, with full training curves shown in Figure 11. VeLO demonstrates competitive performance against Adam with cosine scheduling, while MuLO exhibits significantly degraded performance across the extended training horizon, consistent with the stability patterns observed in vision tasks.
Appendix C Extended Ablation Studies
@@
Vision Transformer Ablations Figures 12 and 13 demonstrate the variation of training dynamics with varying maximum learning rates and weight decay values for the ViT-B/16 model. The results reveal that VeLO exhibits minimal sensitivity to weight decay and scheduler modifications in this experimental setup, suggesting robust inherent regularization properties. Conversely, MuLO shows improved training horizon stability when augmented with cosine scheduling, enabling successful completion of the full 150,000 training steps. Weight decay modifications yield marginal improvements for MuLO across the evaluated parameter ranges.
@@
The language modeling experiments, presented in Figures 14 and 15, demonstrate more pronounced sensitivity to weight decay and scheduling compared to vision tasks. VeLO shows measurable performance improvements when augmented with weight decay and scheduling for specific parameter configurations. Similarly, MuLO benefits substantially from both weight decay regularization and cosine scheduling in the language modeling domain, suggesting that learned optimizers may benefit from scheduling and weight decay.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Appendix D Hyperparameters for Experiments